# A Mutual Information Perspective on Multiple Latent Variable Generative Models for Positive View Generation

### v1

@@ 3.1 Measuring the impact of latent variables in MLVGMs
As described above, for an MLVGM with $n$ latent variables, we optimize the minimax objective of Equation 5 independently for each latent variable $i\in\{1,2,\dots,n\}$. For each $T_{\mathbf{z}_{i}}(\cdot)$, training is stopped when $\mathcal{L}_{\text{InfoNCE}}\approx\overline{\gamma}$, ensuring that the resulting perturbations produce views with a consistent MI shift across all latent variables. This uniformity allows a direct comparison of perturbation magnitudes across different latent spaces.
After training, we obtain $n$ distinct perturbation functions $T_{\mathbf{z}_{i}}(\cdot)$, each tailored to one specific latent variable. Using these functions, we generate image pairs $\mathbf{X}=g(\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,\mathbf{Z}_{i},\dots,\mathbf{Z}_{n};\theta)$ and $\mathbf{X}^{\prime}=g(\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,T_{\mathbf{z}}^{i}(\mathbf{Z}_{i}),\dots,\mathbf{Z}_{n};\theta)$, such that $I(\mathbf{X};\mathbf{X}^{\prime})\approx\gamma$.
More specifically, we perform Monte Carlo (MC) sampling by computing a statistically relevant number of views for each level $i$, estimating the average latent perturbation $\mu_{i}$ required to achieve a similar MI shift in the image space. As depicted in Figure 2, we generally expect that later latent spaces require larger perturbations to achieve the MI shift, matching the empirical observations on the “global-to-local” property of MLVGMs.
In Section 4, we use this strategy to estimate the impact of latent variables for two distinct MLVGMs: a BigBiGan Donahue & Simonyan (2019) pre-trained on ImageNet-$1$K Deng et al. (2009) and a StyleGan2 Karras et al. (2020) pre-trained on LSUN Cars Yu et al. (2015). The former has $6$ latent variables, while the latter has $16$, which we re-organize into $4$ groups of $4$ for computational practicality. These models are subsequently employed for positive view generation, as described in the following.
@@ 3.2 Positive view Generation Strategies
As illustrated in Figure 3(c), we generate pairs of positive views by applying perturbations to one or more latent spaces, each with an appropriately selected magnitude. The choice is guided by the latent impact estimation procedure described above. However, MI shifts alone (or any other metric) can not establish a systematic method for determining optimal positive views. This limitation arises from the intrinsic complexity of the SSCRL problem, which depends on the specific downstream task and the nature of the data itself, as we elaborate in the following.
@@
$(\mathbf{X^{*}_{1}};\mathbf{X^{*}_{2}})=\operatorname*{arg\,min}_{\mathbf{X_{1}};\mathbf{X_{2}}}I(\mathbf{X_{1}};\mathbf{X_{2}});\text{ subject to }I(\mathbf{X_{1}};\mathbf{Y})=I(\mathbf{X_{2}};\mathbf{Y})=I(\mathbf{X};\mathbf{Y});$ (6)
@@
The principle states that optimal views should minimize their Mutual Information (MI) while retaining all information relevant to the downstream task, expressed by some label $\mathbf{Y}$. However, in SSCRL, labels are unavailable, and the downstream task is unknown. Consequently, designing optimal views becomes infeasible. As a result, positive view generation methods focus on obtaining broad applicability across various tasks, relying on heuristics or qualitative evaluation rather than a systematic framework.
To exemplify this phenomenon, we analyze three prominent positive view generation methods. SimCLR (Chen et al., 2020), for instance, defines a broad set of data augmentations $T_{\mathbf{x}}$ to be applied in the pixel space. The specific transformations and their combinations are selected through ablation studies conducted on the ImageNet-$1$K classification task, thereby violating the assumption that the downstream task is unknown.
In the context of Latent Variable Generative Models (LVGMs), two methods for generating views by perturbing the single latent space stand out. First, Jahanian et al. (2021) propose applying random perturbations to an anchor latent variable $\mathbf{z}$. Specifically, the perturbation is defined as $T_{\mathbf{z}}(\mathbf{z})=\mathbf{z}+\mathbf{w}_{\text{rand}}$, where $\mathbf{w}_{\text{rand}}\sim\mathcal{N}^{t}(\mu,\sigma,t)$ follows a truncated Gaussian distribution with truncation parameter $t$. Similar to SimCLR, the parameters of the distribution (e.g., the standard deviation $\sigma$) are tuned via ablation studies, again violating the SSCRL assumption.
In contrast, Li et al. (2022) introduce an adversarial approach to learn the perturbation $T_{\mathbf{z}}$ for each instance. In this case, the positive view is generated as $T_{\mathbf{z}}(\mathbf{z})=\mathbf{z}+\mathbf{w}_{\text{learn}}$, where $\mathbf{w}_{\text{learn}}$ is a learnable perturbation vector. The objective function is formulated similarly to Equation 5, but applied to LVGMs. In this case, the main SSCRL assumption is not violated, as the stopping criterion is decided on empirical inspection of the generated views.
@@
Figure 4: Examples of views generated by adding the same latent vector $\mathbf{w}$ to different levels. (a) Two anchor images and possible views generated by the perturbations of $6$ BigBiGan’s latent levels, represented as the $6$ elements’ vector at the bottom. The darker element indicates the applied perturbation $T_{\mathbf{z}}^{i}(\mathbf{z}_{i})=\mathbf{z}_{i}+\mathbf{w}$ for each latent level $i$. (b) Generated anchors and views by StyleGan2, which has $16$ hierarchical levels, grouped into $4$ sets and represented as the $4$ elements’ vector at the bottom. The darker element indicates the altered group.
In the context of MLVGMs, we maintain the perturbation strategies proposed by Jahanian et al. (2021) and Li et al. (2022), referred to as random and learned, respectively. In fact, our experimental results indicate that the perturbation strategy has a limited impact on the final learned representations. Instead, the magnitude of the perturbations plays a far more critical role. As estimated by our proposed algorithm, MLVGMs offer enhanced control over this aspect, enabling us to tailor magnitudes based on each latent space’s contribution to the generative process.
Similarly to prior approaches, however, defining a fully systematic method for selecting the different magnitudes across $n$ latent spaces remains challenging. While MI helps in quantifying the impact of each latent variable, it does not provide any information on the semantic content of the generated views. To address this limitation, we follow Li et al. (2022) and incorporate qualitative evaluation into the process, ensuring the SSCRL assumption remains intact.
Specifically, we generate multiple examples for each considered MLVGM, by perturbing each latent variable (or group, in the case of StyleGan2) with the same latent vector $\mathbf{w}$, as Figure 4 exemplifies. For BigBigan, trained on a large range of classes, the first latent variable has a significant semantic impact in the generation process, while subsequent variables influence finer details. In contrast, StyleGan2, which operates on the fine-grained domain of vehicles, exhibits a more balanced sub-division. Specifically, the first group of latent variables controls large-scale transformations, such as rotations and zoom. The second group adjusts subject and background composition, while the final groups are primarily responsible for color variations. In summary, when selecting each perturbation magnitude, we consider both these qualitative observations and the quantitative results measured on MI shifts.
@@
The inherent disentanglement of coarse and fine-grained features in MLVGMs offers a clear advantage over standard LVGMs, even when employing similar perturbation strategies (random or learned). To clarify, consider the set of all possible downstream tasks for image data, $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$. For each latent point $\mathbf{z}\sim\mathcal{Z}$, there exists a maximum perturbation magnitude $|\mathbf{w_{z}}|^{\mathcal{T}_{t}}$ for each task $\mathcal{T}_{t}$, such that the resulting views are optimal for the task (Proposition 1).
In SSCRL, however, the downstream task $\mathcal{T}_{t}$ is unknown. Latent perturbation-based positive view generation methods, therefore, aim to define a function $T_{\mathbf{z}}(\mathbf{z})=|\mathbf{w_{z}}|$ that estimates, for each latent point, a non-trivial perturbation that generates valid (true) positives for as many tasks as possible. While larger $|\mathbf{w}|$ can yield more informative (hard) positives, it also increases the likelihood of producing false positives, potentially reducing generalization across diverse tasks.
The primary advantage of MLVGMs lies in their hierarchical structure, which distributes features across multiple latent spaces. This enables the definition of separate perturbation functions $T_{\mathbf{z}}^{i}(\cdot)$ for each latent space $i$, tailored to the impact of that space on the generative process. Crucially, the average perturbation magnitude $|\mathbf{w}|_{i}$ progressively increases with the latent level index $i$, as illustrated in Figure 3(a). For example, perturbations in a high-level latent space (responsible for fine details like textures) can be very large without compromising the validity of positive views for most downstream tasks.
@@ 3.3 Continuous Sampling
With this Continuous Sampling approach, the number of training steps per epoch remains consistent with real-data-based training. However, the total number of unique images seen during training is significantly increased, as the effective training set size becomes $n$ epochs times larger. Moreover, this strategy eliminates the need for pre-generating and storing extensive datasets and avoids standard data-loading bottlenecks, resulting in training times that are comparable to or faster than traditional methods (see Section 4). For a detailed implementation, we provide pseudocode for the continuous sampling procedure in Appendix C.
@@ 4 Experiments
To evaluate the effectiveness of our approach, we compare it against two existing latent perturbation techniques for LVGMs, specifically those introduced by Jahanian et al. (2021) and Li et al. (2022). As an additional baseline, we include SimCLR, a widely-used view generation method applied to real data, and investigate its combination with transformations applied on top of MLVGM-generated views. Furthermore, in Appendix F, we extend the applicability of our method to other generative models beyond GANs by training an NVAE (Vahdat & Kautz, 2020) on the CIFAR-10 dataset (Krizhevsky et al., 2009).
@@ 4.1 Impact of Latent Variables
$1-4$ $0.99$ $15.1$ $2.7$
$5-8$ $1.14$ $29.0$ $4.6$
$9-12$ $0.94$ $38.0$ $5.6$
$13-16$ $0.11$ $134.4$ $14.2$
Following the procedure detailed in Section 3.1, we train $n$ separate perturbation functions $T_{\mathbf{z}}^{i}$ ($n=6$ latent levels for BigBiGan and $n=4$ latent groups for StyleGan2), optimizing the objective in Equation 5. As visually described in Appendix B, the InfoNCE loss rapidly decreases toward zero during the initial training iterations. As the perturbation functions $T_{\mathbf{z}}^{i}$ learn to apply increasing perturbations, the InfoNCE loss rises correspondingly. Training is terminated once a value of $\overline{\gamma}\approx 1$ is achieved, indicating an approximately equal MI shift in the pixel space.
For each latent level or group, we compute the learned perturbation $\mathbf{w_{z}}=p(\mathbf{z};\varphi)$ across a statistically significant number of latent points $\mathbf{z}$. This enables us to estimate the mean ($\mu_{i}$ or $\mu_{g}$) and standard deviation ($\sigma_{i}$ or $\sigma_{g}$) of the inferred perturbation distributions $q^{i}(|\mathbf{w}|)$ or $q^{g}(|\mathbf{w}|)$. Table 1 presents these results, along with the final InfoNCE loss achieved during training.
From Table 1(a) (Monte Carlo results for BigBiGan), we observe that the average perturbation (estimated mean $\mu_{i}$) required to achieve a comparable InfoNCE loss increases progressively across latent levels, from $i=1$ to $i=5$. However, for $i=6$, the InfoNCE loss does not rise substantially even under high average perturbations, suggesting an under-utilization of the latent level in the generative process. Conversely, we measure a very low $\mu_{i}$ for the first latent level, suggesting a possible over-utilization. These observations may indicate potential inefficiencies in the BigBiGan architecture or training procedure.
A similar trend is observed for StyleGan2 (Table 1(b)), where larger perturbation magnitudes ($\mu_{g}$) are needed to achieve comparable InfoNCE loss values as latent groups progress from $g=1-4$ to $g=13-16$. Notably, the final group exhibits a degenerate behavior, where even large perturbations fail to influence the MI of the generated views significantly.
@@ 4.2 View Generation
Table 2: Comparison of representations learned on the ImageNet-$1$K dataset or BigBiGan generator with two contrastive frameworks (SimCLR and SimSiam). Metrics are Top-$1$ and Top-$5$ accuracy for linear classification on ImageNet-$1$K, average precision for detection on Pascal VOC, and mean Top-$1$ accuracy over $7$ transfer classification datasets. Bold indicates the best result for each group, underline the absolute best, and indicates the baseline reported from Li et al. (2022).
@@
ImageNet-$1$K Pascal VOC ImageNet-$1$K Pascal VOC Transfer
Top-$1$ Top-$5$ AP $\text{AP}_{50}$ $\text{AP}_{75}$ Top-$1$ Top-$5$ AP $\text{AP}_{50}$ $\text{AP}_{75}$ Top-$1$
@@
Table 3: Comparison of representations learned on the LSUN Cars dataset or StyleGan 2 generator with two contrastive frameworks (SimSiam and Byol). Metrics are Top-$1$ and Top-$5$ accuracy for linear classification on Stanford Cars and FGVC Aircraft. Bold indicates the best result for each group, underline the absolute best.
@@
We test MLVGMs generated views by training multiple ResNet-$50$ encoders, using SimSiam (Chen & He, 2021), SimCLR (Chen et al., 2020) (on BigBiGan, following previous work Li et al. (2022)) and Byol (Grill et al., 2020) (on StyleGan2). We sample latent anchors from a truncated normal distribution: $\mathcal{N}^{t}(0.0,1.0,2.0)$ for BigBiGan and $\mathcal{N}^{t}(0.0,1.0,0.9)$ for StyleGan2. Positives are computed using the random or learned strategies, applied separately on each latent level. Given the views visualization of Figure 3 and the MC results reported in Table 1, we fix the first latent level on BigBiGan and apply only tiny perturbations on the first two groups of StyleGan2, in order to not affect sensible semantic aspects. Conversely, we enhance the perturbations of the remaining levels, which alter more local details/colors. The specific magnitudes, as well as other hyperparameters, are reported in Appendix D.
@@
The representation capabilities of the obtained encoders are compared against several methods: training on synthetic data without latent perturbations $T_{\mathbf{z}}$, the random and learned baselines using single latent spaces, and the upper bound of using real data ($1.28$M images for ImageNet-$1$K (Deng et al., 2009) and $893$K images for LSUN Cars (Yu et al., 2015)). In all these scenarios, SimCLR pixel-space augmentations $T_{\mathbf{x}}$ are used, consisting of random cropping, horizontal flipping, grayscale, and color jittering. Since our ML views generate realistic color changes (see Appendix H), we only partially apply $T_{\mathbf{x}}$ transformations on top of our positives, removing grayscale and color jittering. To better investigate this aspect, in Appendix E we further test various combinations of $T_{\mathbf{x}}$ coupled with our method.
@@
BigBiGan views are evaluated on ImageNet-$1$K linear classification and, for Simsiam, on seven transfer datasets: Birdsnap (Berg et al., 2014), Caltech101101101101 (Fei-Fei et al., 2004), Cifar100100100100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), Flowers102102102102 (Nilsback & Zisserman, 2008), Food101101101101 (Bossard et al., 2014), and Pets (Parkhi et al., 2012). We also compute Average Precision on Pascal VOC (Everingham et al., 2010) object detection using detectron 2 (Wu et al., 2019) to train a Faster-RCNN with the R50505050-C4444 backbone. The results are reported in Table 2, including the mean accuracy for the transfer tasks (complete runs in Appendix G). For StyleGan2, we compute linear classification accuracy on Stanford Cars (Krause et al., 2013) and FGCV Aircraft $2013$b (Maji et al., 2013), reporting results in Table 3.
@@
In all experiments, MLVGMs views outperform the corresponding baseline, proving their superior quality. Comparing random and learned methods, we observe that the multiple latent (ML) random experiments often close the gap with the learned counterparts. This suggests that distinct-level perturbations are more important than the selected alteration technique. In comparison with real data, ML views generally yield better or similar results, except in the case of SimSiam encoders evaluated on ImageNet-$1$K. However, this gap narrows or disappears in other downstream tasks and datasets, evidencing good generalization capabilities of the learned representations, which is the main goal of SSCRL. For StyleGan2, the great performance boost given by ML random views on Stanford Cars is noteworthy. When generalizing to FGCV Aircraft, all runs achieve similar performance, with marginal improvements of the ML runs when using SimSiam, and good real data results on Byol. This may be due to the high domain shift between the two datasets (Car vs Aircraft), leading to a challenging generalization for all representations.

### v2

@@ 3.1 Measuring the impact of latent variables in MLVGMs
As a result of the above, we obtain $n$ independent perturbation functions $T_{\mathbf{z}_{i}}(\cdot)$, acting on single latents. For each of these, training is stopped when $\mathcal{L}_{\text{InfoNCE}}\approx\overline{\gamma}$, ensuring views with a consistent MI shift across all latent variables, and allowing the direct comparison of perturbation magnitudes across different latent spaces.
Afterwards, we perform Monte Carlo (MC) sampling by computing a statistically relevant number of image pairs for each level $i$: $\mathbf{X}=g(\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,\mathbf{Z}_{i},\dots,\mathbf{Z}_{n};\theta)$ and $\mathbf{X}^{\prime}=g(\mathbf{Z}_{1},\mathbf{Z}_{2},\dots,T_{\mathbf{z}}^{i}(\mathbf{Z}_{i}),\dots,\mathbf{Z}_{n};\theta)$, where we know that $I(\mathbf{X};\mathbf{X}^{\prime})\approx\overline{\gamma}$. This enables to estimate the average latent perturbation $\mu_{i}$ required to achieve a similar MI shift in the image space. As depicted in Figure˜2, we generally expect that later latent spaces require larger perturbations to achieve the MI shift, matching the empirical observations on the “global-to-local” property of MLVGMs.
In Section˜4, we use this strategy to estimate the impact of latent variables for two distinct MLVGMs: a BigBiGan Donahue & Simonyan (2019) pre-trained on ImageNet-11K Deng et al. (2009) and a StyleGan2 Karras et al. (2020) pre-trained on LSUN Cars Yu et al. (2015). The former has $6$ latent variables, while the latter has $16$, which we re-organize into $4$ groups of $4$ for computational practicality.
@@ 3.2 Positive view Generation Strategies
As illustrated in Figure˜3(c), we generate pairs of positive views by applying perturbations to one or more latent spaces, each with an appropriately selected magnitude. To better understand the rationale behind our approach, we first look at how previous methods decide view generation strategies.
@@
[…]
@@
The principle states that optimal views should minimize their Mutual Information (MI) while retaining all information relevant to the downstream task, expressed by some label $\mathbf{Y}$. However, in SSCRL, labels are unavailable, and the downstream task is unknown. Consequently, designing optimal views becomes infeasible. To solve this inherent shortcoming, most methods design views by implicitly fixing an MI threshold that decides true positives:
Definition 2
(Mutual Information bound for SSCRL).
Let $\mathbf{X}_{1},\mathbf{X}_{2}$ be two random variables in the common image space. In absence of a known downstream task $\mathcal{T}$ with labels $\mathbf{Y}\in\mathcal{Y}$; the variables $\mathbf{X}_{1},\mathbf{X}_{2}$ can be considered as positive views if
$I\bigl(\mathbf{X}_{1},\mathbf{X}_{2}\bigr)\geq\rho;$ (7)
where $\rho$ is an implicitly defined MI threshold.
In other terms, the goal of view generation is to define positives that share a reasonable amount of information, relevant for as many tasks as possible. To exemplify this phenomenon, we analyze three prominent positive view generation methods. SimCLR (Chen et al., 2020), defines a broad set of data augmentations $T_{\mathbf{x}}$ to be applied in the pixel space. The specific transformations and their combinations are selected through ablation studies conducted on the ImageNet-11K classification task. In this context, the threshold $\rho$ is implicitly fixed by means of Equation˜6, where a pretext downstream task is considered.
In the context of Latent Variable Generative Models (LVGMs), two methods for generating views by perturbing the single latent space stand out. First, Jahanian et al. (2021) propose applying random perturbations to an anchor latent variable $\mathbf{z}$. Specifically, the perturbation is defined as $T_{\mathbf{z}}(\mathbf{z})=\mathbf{z}+\mathbf{w}_{\text{rand}}$, where $\mathbf{w}_{\text{rand}}\sim\mathcal{N}^{t}(\mu,\sigma,t)$ follows a truncated Gaussian distribution with truncation parameter $t$. Similar to SimCLR, the parameters of the distribution (e.g., the standard deviation $\sigma$) are tuned via ablation studies on ImageNet, fixing $\rho$ with a pretext task.
In contrast, Li et al. (2022) introduce an adversarial approach to learn the perturbation $T_{\mathbf{z}}$ for each instance. In this case, the positive view is generated as $T_{\mathbf{z}}(\mathbf{z})=\mathbf{z}+\mathbf{w}_{\text{learn}}$, where $\mathbf{w}_{\text{learn}}$ is a learnable perturbation vector. The objective function is formulated similarly to Equation˜5, but applied to LVGMs. In this case, the training stopping criterion (and therefore the threshold $\rho$) is empirically decided by observing the quality of generated views at each step.
@@
We maintain the perturbation strategies proposed by Jahanian et al. (2021) and Li et al. (2022), referred to as random and learned, respectively. In fact, the core advantage of our method relies on designing tailored magnitudes based on each latent space’s contribution to the generative process, rather than proposing a novel perturbation strategy.
To do so, we proceed with two distinct phases. First, we use the quantification algorithm to relate the impact of different latents. This reveals the presence of over- or under-used codes, which we can immediately discard. Overused codes imply that even very small perturbations will result in a low $\rho$ threshold, while unused codes do not affect the threshold in any way. Therefore, this first step has the goal to discard perturbations that lead to non-informative views. Finally, in the second phase we consider the remaining codes only and visualize the semantic effect of the applied perturbations to estimate valid magnitudes, implicitly defining $\rho$.
Figure 4: Examples of views generated by adding the same latent vector $\mathbf{w}$ to different levels. (a) Two anchor images and possible views generated by the perturbations of 66 BigBiGan’s latent levels, represented as the 66 elements’ vector at the bottom. The darker element indicates the applied perturbation $T_{\mathbf{z}}^{i}(\mathbf{z}_{i})=\mathbf{z}_{i}+\mathbf{w}$ for each latent level ii. (b) Generated anchors and views by StyleGan2, which has $16$ hierarchical levels, grouped into 44 sets and represented as the 44 elements’ vector at the bottom. The darker element indicates the altered group.
As an example, in Figure˜4 we generate multiple examples for each considered MLVGM, by perturbing each latent variable (or group, in the case of StyleGan2) with the same latent vector $\mathbf{w}$. For BigBiGan, the quantitative analysis reveals that the first code is overused, and we can verify that even small perturbations completely change the views semantic content. On the opposite, the last latent is underused, and visually presents no differences w.r.t. the reference image. As a result, we apply perturbations only on the remaining four latents, which act on small shape details or colors. Similarly, for StyleGan2, we first quantitatively verify that the last group is underused, and discard it. For the remaining ones, we visually observe how the first group controls large-scale transformations, the second group adjusts subject and background composition, while the latter is primarily responsible for color variations.
@@
The inherent disentanglement of coarse and fine-grained features in MLVGMs offers a clear advantage over standard LVGMs, even when employing similar perturbation strategies (random or learned). To clarify, consider the set of all possible downstream tasks for image data, $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$. For each latent point $\mathbf{z}\sim\mathcal{Z}$, there exists a maximum perturbation magnitude $|\mathbf{w_{z}}|^{\mathcal{T}_{t}}$ for each task $\mathcal{T}_{t}$, such that the resulting views are optimal for the task (Proposition˜1).
In SSCRL, however, the downstream task $\mathcal{T}_{t}$ is unknown. Latent perturbation-based positive view generation methods, therefore, aim to define a function $T_{\mathbf{z}}(\mathbf{z})=|\mathbf{w_{z}}|$ that estimates, for each latent point, a non-trivial perturbation that generates valid (true) positives for as many tasks as possible. In other terms, it exists a direct mapping between the selected latent functions and the choice of the MI threshold $\rho$: while larger $|\mathbf{w}|$ (lower $\rho$) can yield more informative (hard) positives, it also increases the likelihood of producing false positives, potentially reducing generalization across diverse tasks.
The primary advantage of MLVGMs lies in their hierarchical structure, which distributes features across multiple latent spaces. This enables the definition of separate perturbation functions $T_{\mathbf{z}}^{i}(\cdot)$ for each latent space $i$, tailored to the impact of that space on the generative process. Crucially, the average perturbation magnitude $|\mathbf{w}|_{i}$ progressively increases with the latent level index $i$, as illustrated in Figure˜3(a). For example, perturbations in a high-level latent space (responsible for fine details like textures) can be very large without compromising the validity of positive views for most downstream tasks - i.e. they have a low effect on $\rho$.
@@ 3.3 Continuous Sampling
With this Continuous Sampling approach, the number of training steps per epoch remains consistent with real-data-based training. However, the total number of unique images seen during training is significantly increased, as the effective training set size becomes $n$ epochs times larger. Moreover, this strategy eliminates the need for pre-generating and storing extensive datasets and avoids standard data-loading bottlenecks, resulting in training times that are comparable to or faster than traditional methods (see Section˜4). For a detailed implementation, we provide pseudocode for the continuous sampling procedure in Appendix˜C.
@@ 4 Experiments
To evaluate the effectiveness of our approach, we compare it against two existing latent perturbation techniques for LVGMs, specifically those introduced by Jahanian et al. (2021) and Li et al. (2022). As an additional baseline, we include SimCLR, a widely-used view generation method applied to real data, and investigate its combination with transformations applied on top of MLVGM-generated views. Furthermore, in Appendix˜F, we extend the applicability of our method to other generative models beyond GANs by training an NVAE (Vahdat & Kautz, 2020) on the CIFAR-10 dataset (Krizhevsky et al., 2009).
@@ 4.1 Impact of Latent Variables
141-4 $0.99$ $15.1$ $2.7$
585-8 $1.14$ $29.0$ $4.6$
9129-12 $0.94$ $38.0$ $5.6$
131613-16 $0.11$ $134.4$ $14.2$
Following the procedure detailed in Section˜3.1, we train $n$ separate perturbation functions $T_{\mathbf{z}}^{i}$ ($n=6$ latent levels for BigBiGan and $n=4$ latent groups for StyleGan2), optimizing the objective in Equation˜5. As visually described in Appendix˜B, the InfoNCE loss rapidly decreases toward zero during the initial training iterations. As the perturbation functions $T_{\mathbf{z}}^{i}$ learn to apply increasing perturbations, the InfoNCE loss rises correspondingly. Training is terminated once a value of $\overline{\gamma}\approx 1$ is achieved, indicating an approximately equal MI shift in the pixel space.
For each latent level or group, we compute the learned perturbation $\mathbf{w_{z}}=p(\mathbf{z};\varphi)$ across a statistically significant number of latent points $\mathbf{z}$. This enables us to estimate the mean ($\mu_{i}$ or $\mu_{g}$) and standard deviation ($\sigma_{i}$ or $\sigma_{g}$) of the inferred perturbation distributions $q^{i}(|\mathbf{w}|)$ or $q^{g}(|\mathbf{w}|)$. Table˜1 presents these results, along with the final InfoNCE loss achieved during training.
From Table˜1(a) (Monte Carlo results for BigBiGan), we observe that the average perturbation (estimated mean $\mu_{i}$) required to achieve a comparable InfoNCE loss increases progressively across latent levels, from $i=1$ to $i=5$. However, for $i=6$, the InfoNCE loss does not rise substantially even under high average perturbations, suggesting an under-utilization of the latent level in the generative process. Conversely, we measure a very low $\mu_{i}$ for the first latent level, suggesting a possible over-utilization. These observations may indicate potential inefficiencies in the BigBiGan architecture or training procedure.
A similar trend is observed for StyleGan2 (Table˜1(b)), where larger perturbation magnitudes ($\mu_{g}$) are needed to achieve comparable InfoNCE loss values as latent groups progress from g=14g=1-4 to g=1316g=13-16. Notably, the final group exhibits a degenerate behavior, where even large perturbations fail to influence the MI of the generated views significantly.
@@ 4.2 View Generation
Table 2: Comparison of representations learned on the ImageNet-11K dataset or BigBiGan generator with two contrastive frameworks (SimCLR and SimSiam). Metrics are Top-11 and Top-$5$ accuracy for linear classification on ImageNet-11K, average precision for detection on Pascal VOC, and mean Top-$1$ accuracy over 77 transfer classification datasets. “random” row refers to Jahanian et al. (2021), and “learned” to Li et al. (2022). Bold indicates the best result for each group, underline the absolute best, and indicates the baseline reported from Li et al. (2022).
@@
ImageNet-11K Pascal VOC ImageNet-11K Pascal VOC Transfer
Top-11 Top-55 AP $\text{AP}_{50}$ $\text{AP}_{75}$ Top-11 Top-55 AP $\text{AP}_{50}$ $\text{AP}_{75}$ Top-11
@@
Table 3: Comparison of representations learned on the LSUN Cars dataset or StyleGan 2 generator with two contrastive frameworks (SimSiam and Byol). Metrics are Top-11 and Top-$5$ accuracy for linear classification on Stanford Cars and FGVC Aircraft. “random” row refers to Jahanian et al. (2021), and “learned” to Li et al. (2022). Bold indicates the best result for each group, underline the absolute best.
@@
We test MLVGMs generated views by training multiple ResNet-$50$ encoders, using SimSiam (Chen & He, 2021), SimCLR (Chen et al., 2020) (on BigBiGan, following previous work Li et al. (2022)) and Byol (Grill et al., 2020) (on StyleGan2). We sample latent anchors from a truncated normal distribution: $\mathcal{N}^{t}(0.0,1.0,2.0)$ for BigBiGan and $\mathcal{N}^{t}(0.0,1.0,0.9)$ for StyleGan2. Positives are computed using the random or learned strategies, applied separately on each latent level. Given the MC results reported in Table˜1, we first select the latents to discard. These are the first latent of BigBiGan (overused) and the last latent/group in both BigBiGan and StyleGan2 (underused). Then, we systematically apply different perturbations to the remaining latents, observing that latents 252-5 in BigBiGan obtain relevant semantic changes with a similar perturbation magnitude. For StyleGan2, we apply only tiny perturbations on the first two groups (modify shape and global transformations), and a larger perturbation on the remaining group (alters colors). The specific magnitudes, as well as other hyperparameters, are reported in Appendix˜D.
@@
The representation capabilities of the obtained encoders are compared against several methods: training on synthetic data without latent perturbations $T_{\mathbf{z}}$, the random and learned baselines using single latent spaces, and the upper bound of using real data (1.281.28M images for ImageNet-11K (Deng et al., 2009) and 893893K images for LSUN Cars (Yu et al., 2015)). In all these scenarios, SimCLR pixel-space augmentations $T_{\mathbf{x}}$ are used, consisting of random cropping, horizontal flipping, grayscale, and color jittering. Since our ML views generate realistic color changes (see Appendix˜H), we only partially apply $T_{\mathbf{x}}$ transformations on top of our positives, removing grayscale and color jittering. To better investigate this aspect, in Appendix˜E we further test various combinations of $T_{\mathbf{x}}$ coupled with our method.
@@
BigBiGan views are evaluated on ImageNet-11K linear classification and, for Simsiam, on seven transfer datasets: Birdsnap (Berg et al., 2014), Caltech101101 (Fei-Fei et al., 2004), Cifar100100 (Krizhevsky et al., 2009), DTD (Cimpoi et al., 2014), Flowers102102 (Nilsback & Zisserman, 2008), Food101101 (Bossard et al., 2014), and Pets (Parkhi et al., 2012). We also compute Average Precision on Pascal VOC (Everingham et al., 2010) object detection using detectron 2 (Wu et al., 2019) to train a Faster-RCNN with the R5050-C44 backbone. The results are reported in Table˜2, including the mean accuracy for the transfer tasks (complete runs in Appendix˜G). For StyleGan2, we compute linear classification accuracy on Stanford Cars (Krause et al., 2013) and FGCV Aircraft 20132013b (Maji et al., 2013), reporting results in Table˜3.
@@
In all experiments, MLVGMs views outperform the corresponding baseline, proving their superior quality. Comparing random and learned methods, we observe that the multiple latent (ML) random experiments often close the gap with the learned counterparts. This suggests that distinct-level perturbations are more important than the selected alteration technique. In comparison with real data, ML views generally yield better or similar results, except in the case of SimSiam encoders evaluated on ImageNet-11K. However, this gap narrows or disappears in other downstream tasks and datasets, evidencing good generalization capabilities of the learned representations, which is the main goal of SSCRL. For StyleGan2, the great performance boost given by ML random views on Stanford Cars is noteworthy. When generalizing to FGCV Aircraft, all runs achieve similar performance, with marginal improvements of the ML runs when using SimSiam, and good real data results on Byol. This may be due to the high domain shift between the two datasets (Car vs Aircraft), leading to a challenging generalization for all representations.